# Qubit-assisted quantum metrology

### v1

@@ VI Discussion
where the coefficients of the phase generator are
$c_{z}$ $=-\frac{g^{2}}{\tilde{\omega}^{2}}\cos(\tilde{\omega}t_{1})-\frac{\omega_{P}^{2}}{\tilde{\omega}^{2}},$ (38)
$c_{x}$ $=\frac{g\omega_{P}}{\tilde{\omega}^{2}}[\cos(\tilde{\omega}t_{1})-1],$
$c_{y}$ $=-\frac{g}{\tilde{\omega}}\sin(\tilde{\omega}t_{1}).$
We are now on the stage of minimizing the magnitude of the second term in Eq. (16) to enhance QFI. With Eqs. (37), (14), and (15), we have
$\langle\psi_{i}|\langle\varphi|U_{\theta}^{\dagger}(\partial_{\theta}U_{\theta})|\varphi\rangle|\psi_{i}\rangle$ (39)
$=$ $c_{z}\langle J_{z}\rangle_{\psi_{i}}+\langle\sigma_{z}\rangle_{\varphi}\langle c_{x}J_{x}+c_{y}J_{y}\rangle_{\psi_{i}}.$
@@
$\theta_{0}=\frac{\pi}{2},\quad t_{1}=\tilde{\omega}^{-1}\left[\arccos\left(-\frac{\omega_{P}^{2}}{g^{2}}\right)+2n\pi\right],$ (40)
with an integer $n$, both terms in Eq. (39) can vanish at the same time for arbitrary probe state $\psi_{i}$.
Using Eqs. (16) and (40), QFI becomes the mean square of a phase generator, i.e.,
$F_{Q}=4\sum_{i=1}^{d}p_{i}\left\langle(c_{x}J_{x}+c_{y}J_{y})^{2}\right\rangle_{\psi_{i}}.$ (41)
It reaches the maximum value $F_{Q}=N^{2}$ when the probe spin ensemble is a pure state $|j,j\rangle^{\prime}_{\rm opt}$, or $|j,-j\rangle^{\prime}_{\rm opt}$, and or a mixed state over them. Here $|j,m\rangle^{\prime}_{\rm opt}$ with $-j\leq m\leq j$ denotes the eigenstate for the optimized collective angular momentum operator $c_{x}J_{x}+c_{y}J_{y}$.
@@
Since $\arccos(x)$ ranges from $-1$ to $1$, the optimal evolution time $t_{1}$ in Eq. (40) can only be obtained in the deep strong coupling regime $g\geq\omega_{P}$. When $g<\omega_{P}$, only the second term in Eq. (39) can vanish for arbitrary probe state and the minimum magnitude of the first term associated with $c_{z}$ can be found at $\cos(\tilde{\omega}t_{1})=-1$. The optimal condition, that does not render $F_{Q}=N^{2}$, can thus be rewritten as
$\theta_{0}=\frac{\pi}{2},\quad t_{1}=\frac{(2n+1)\pi}{\tilde{\omega}},$ (42)
@@
Figure 5: QFI as a function of $N$ for $t_{1}=\pi/(\sqrt{2}g)$ and $g/\omega_{P}=1$ (red crosses); for $t_{1}=\pi/(3.48g)$ and $g/\omega_{P}=0.3$ (blue circles); for $t_{1}=\pi/(\sqrt{101}g)$ and $g/\omega_{P}=0.1$ (orange squares). The orange solid line is a fit curve for the last case with $F_{Q}=aN^{2}+bN$, where $a\approx 0.04$ and $b\approx 0.96$. The black dashed line and the black dot-dashed lines indicate the Heisenberg scaling $N^{2}$ and the shot-noise scaling $N$.
@@
In Fig. 5, we show the dependence of QFI on the probe-spin number $N$ under various $t_{1}$ and $g/\omega_{P}$ according to Eq. (40) or Eq. (42) with $n=0$, where the initial state of the probe ensemble is set as $\rho_{P}^{\prime\rm opt}=|j,j\rangle^{\prime}_{\rm opt}\langle j,j|$. Numerical simulation confirms that when $g=\omega_{P}$, QFI attains its peak value $F_{Q}=N^{2}$. The scaling behavior of QFI gradually deviates from the Heisenberg limit with a decreasing $g$. Yet the numerical fitting for $g/\omega_{P}=0.1$ (see the orange solid line and the orange squares) indicate that QFI will follow the Heisenberg scaling in an asymptotic way. It is found that $F_{Q}\approx 0.04N^{2}+0.96N$, where the two coefficients are dependent on $g/\omega_{P}$ yet independent of $N$.
VII Conclusion
@@
In summary, we incorporate the joint-evolution-and-quantum-measurement idea to quantum parameter estimation in our metrology protocol. With no input of entangled or squeezed state and with no use of nonlinear Hamiltonian, QFI about the phase encoded into the probe system can be optimized to be the mean square of an optimized phase generator with respect to the probe initial state. That renders an exact or asymptotic Heisenberg-scaling behavior in terms of the spin number $N$ even when the probe starts from a thermal state. And the numerical simulation shows that the imperfections in preparing the ancillary qubit and optimizing the evolution time does not change the Heisenberg-scaling behavior about metrology precision as long as $N$ is sufficiently large. By virtue of the time-reversal strategy and the projective measurement on the whole system or merely the ancillary qubit, CFI is found to be saturated with its quantum counterpart. Our protocol of qubit-assisted quantum metrology is applicable to a general spin-spin-bath model. It therefore provides an economical way toward the Heisenberg-scaling metrology.
@@ Acknowledgments
Appendix A Evolution operator for time-reversal strategy in the qubit-assisted metrology
This appendix derives the unitary evolution operator for the time-reversal strategy in our qubit-assisted metrology. With the time-reversal condition (3), the whole evolution operator in Eq. (12) can be written as
@@ Appendix A Evolution operator for time-reversal strategy in the qubit-assisted metrology
In the last step, we have used $e^{i\alpha J_{z}}J_{x}e^{-i\alpha J_{z}}=\cos(\alpha)J_{x}-\sin(\alpha)J_{y}$. The Baker-Campbell-Hausdorff formula yields:
@@
where $C_{k+1}=igt_{1}[J_{z}\sigma_{z},C_{k}]$ and $C_{0}=J_{x}$. Using the commutation relation $[J_{\alpha},J_{\beta}]=i\epsilon_{\alpha\beta\gamma}J_{\gamma}$ with $\epsilon_{\alpha\beta\gamma}$ the Levi-Civita tensor, we have

### v2

@@ V Classical Fisher information
where the coefficients of the phase generators are
$c_{z}$ $=-\frac{g^{2}}{\tilde{\omega}^{2}}\cos(\tilde{\omega}t_{1})-\frac{\omega_{P}^{2}}{\tilde{\omega}^{2}},$ (40a)
$c_{x}$ $=\frac{g\omega_{P}}{\tilde{\omega}^{2}}[\cos(\tilde{\omega}t_{1})-1],$ (40b)
$c_{y}$ $=-\frac{g}{\tilde{\omega}}\sin(\tilde{\omega}t_{1}).$ (40c)
We are now on the stage of minimizing the magnitude of the second term in Eq. (16) to ensure the quadratic scaling of QFI. With Eqs. (39), (14), and (15), we have
$\langle\psi_{i}|\langle\varphi|U_{\theta}^{\dagger}(\partial_{\theta}U_{\theta})|\varphi\rangle|\psi_{j}\rangle$ (41)
$=$ $ic_{z}\langle\psi_{i}|J_{z}|\psi_{j}\rangle+i\langle\sigma_{z}\rangle_{\varphi}\langle\psi_{i}|c_{x}J_{x}+c_{y}J_{y}|\psi_{j}\rangle.$
@@
$\theta_{0}=\frac{\pi}{2},\quad t_{1}=\tilde{\omega}^{-1}\left[\arccos\left(-\frac{\omega_{P}^{2}}{g^{2}}\right)+2n\pi\right]$ (42)
with an integer $n$, the two terms in Eq. (41) can vanish at the same time for arbitrary probe state $|\psi_{i}\rangle$.
By using Eqs. (16) and (42), QFI becomes the mean square of a phase generator, i.e.,
$F_{Q}=4\sum_{i=1}^{d}p_{i}\left\langle\left(c_{x}J_{x}+c_{y}J_{y}\right)^{2}\right\rangle_{\psi_{i}}.$ (43)
It reaches the maximum value $F_{Q}=N^{2}$ when the probe spin ensemble is in a pure state $|j,j\rangle^{\prime}_{\rm opt}$, $|j,-j\rangle^{\prime}_{\rm opt}$, or an arbitrary mixed or superposed state over them. Here $|j,m\rangle^{\prime}_{\rm opt}$ with $-j\leq m\leq j$ denotes the eigenstate of the optimized phase generator $c_{x}J_{x}+c_{y}J_{y}$.
Figure 5: QFI as a function of $N$ for $gt_{1}=\pi/\sqrt{2}$ and $g/\omega_{P}=1$ (red crosses); for $gt_{1}=\pi/3.48$ and $g/\omega_{P}=0.3$ (blue circles); for $gt_{1}=\pi/\sqrt{101}$ and $g/\omega_{P}=0.1$ (orange squares). The orange solid line is a fit curve for the last case with $F_{Q}=aN^{2}+bN$, where $a\approx 0.04$ and $b\approx 0.96$. The black dashed line and the black dot-dashed line indicate the Heisenberg scaling and the shot-noise scaling, respectively.
@@
Since $\arccos(x)$ is defined for $x$ from $-1$ to $1$, the optimal evolution time $t_{1}$ in Eq. (42) can only be accessible in the deep strong coupling regime $g\geq\omega_{P}$. When $g<\omega_{P}$, only the second term in Eq. (41) can vanish for arbitrary probe state and the minimum magnitude of the first term associated with $c_{z}$ can be found at $\cos(\tilde{\omega}t_{1})=-1$ due to Eq. (40a). The sub-optimized condition, that does not render $F_{Q}=N^{2}$, can thus be rewritten as
$\theta_{0}=\frac{\pi}{2},\quad t_{1}=\frac{(2n+1)\pi}{\tilde{\omega}}$ (44)
@@
In Fig. 5, we show the dependence of QFI on the probe-spin number $N$ under various $g/\omega_{P}$, where $t_{1}$ is specified by Eq. (42) or Eq. (44) with $n=0$ and the probe ensemble is initialized as $\rho_{P}^{\prime\rm opt}=|j,j\rangle^{\prime}_{\rm opt}\langle j,j|$. Numerical simulation confirms that when $g=\omega_{P}$, QFI attains its peak value $F_{Q}=N^{2}$ (see the red crosses). The scaling behavior of QFI gradually deviates from the Heisenberg limit with a decreasing $g$. The evolution time $t_{1}$ and the data for QFI presented by the blue circles and the orange squares are obtained with the sub-optimized condition in Eq. (44). However, the numerical simulation and fitting for $g/\omega_{P}=0.1$ (see the orange solid line and the orange squares) indicate that QFI still follow the Heisenberg scaling in an asymptotic way. It is found that $F_{Q}\approx 0.04N^{2}+0.96N$, where the two factors are dependent on the ratio $g/\omega_{P}$ and independent of $N$.
VII Discussion and Conclusion
Decoherence can damage any quantum metrology protocol. Here we incorporate the dephasing noise on the ancillary qubit to the ideal model under the Hamiltonian (1). The noise can be described by the following two Kraus operators Demkowicz-Dobrzański and Maccone (2014):
$K_{1}=\sqrt{1-\frac{x}{2}}\mathcal{I}^{2},\quad K_{2}=\sqrt{\frac{x}{2}}\sigma_{z},$ (45)
where $x\in[0,1]$ indicates the decoherence rate. The noise-free situation can be recovered when $x=0$. Since the two operators commute with the parametric-encoded unitary evolution operator $U_{\theta}$ as given in Eq. (13), they can be directly inserted into the entire evolution. We then have
$\rho(\theta)$ $=U_{\theta}\rho_{P}\otimes\left(\sum_{k=1,2}K_{k}\rho_{A}K_{k}^{\dagger}\right)U_{\theta}^{\dagger}$ (46)
$=U_{\theta}\rho_{P}\otimes\rho_{A}^{\prime}U_{\theta}^{\dagger}.$
Under the optimal condition in Eq. (18), QFI of the output state becomes
$F_{Q}$ $=4\sum_{i=1}^{d}p_{i}\langle J_{\rm opt}^{2}\rangle_{\psi_{i}}$ (47)
$-\sum_{i,j=1}^{d}\frac{8x(2-x)p_{i}p_{j}}{(2-x)p_{i}+xp_{j}}|\langle\psi_{i}|J_{\rm opt}|\psi_{j}\rangle|^{2}.$
It is found that the ideal result in Eq. (21) can be immediately restored with $x=0$. Under an optimized state $|j,j\rangle_{\rm opt}$ for the probe system in the noise-free situation, we have
$F_{Q}=(1-x)^{2}N^{2}.$ (48)
It turns out that QFI is reduced, yet still follows the Heisenberg scaling behavior $F\propto N^{2}$ under the dephasing noise.
@@ VII Discussion and Conclusion
In summary, we incorporate the joint-evolution-and-quantum-measurement idea to the parameter estimation in our quantum metrology protocol, by which the quantum Fisher information for the phase encoded in the probe system (a spin ensemble) can be optimized to be the mean square of a parametric-dependent phase generator with respect to the probe initial state. It is a metrology protocol for atomic systems with no squeezed or GHZ state and with no nonlinear Hamiltonian. That renders an exact or asymptotic Heisenberg-scaling behavior in the size of the probe system $N$ even when the probe starts from a thermal state. Our calculation shows that the Heisenberg-scaling behavior about metrology precision holds under imperfect controls over the initial state of the ancillary qubit, the optimized evolution time, and the system parameters as long as $N$ is sufficiently large. It even demonstrates certain robustness against the dephasing noise. By virtue of the time-reversal strategy and the projective measurement on the entire system or simply on the ancillary qubit, the classical Fisher information is found to be saturated with its quantum counterpart. Our protocol is applicable to a general spin-spin-bath model, providing insights for enhancing quantum metrology by quantum measurement. In essence, it paves an economical way toward the Heisenberg-scaling metrology.
@@ Acknowledgments
Appendix A Evolution operator for qubit-assisted circuit under time-reversal strategy
This appendix derives the unitary evolution operator for the qubit-assisted circuit model in Fig. 1. With the time-reversal condition (3), the entire evolution operator in Eq. (12) can be derived as
@@ Appendix A Evolution operator for qubit-assisted circuit under time-reversal strategy
$=e^{iHt_{1}}e^{-iHt_{1}}+\left(-i\theta e^{iHt_{1}}J_{x}e^{-iHt_{1}}\right)$
$+\frac{1}{2!}\left(-i\theta e^{iHt_{1}}J_{x}\right)\left(-i\theta J_{x}e^{-iHt_{1}}\right)+\cdots$
$=\mathcal{I}+\left(-i\theta e^{iHt_{1}}J_{x}e^{-iHt_{1}}\right)$
$+\frac{1}{2!}\left(-i\theta e^{iHt_{1}}J_{x}e^{-iHt_{1}}\right)\left(-i\theta e^{iHt_{1}}J_{x}e^{-iHt_{1}}\right)+\cdots$
$+\frac{1}{k!}\left(-i\theta e^{iHt_{1}}J_{x}e^{-iHt_{1}}\right)^{k}+\cdots$
@@
in which the operators $e^{\pm iHt_{1}}$ are put inside the power expansion of the exponential $e^{-i\theta J_{x}}$.
@@
In the first equivalence, we employ the identity that
$e^{i\alpha J_{z}}J_{x}e^{-i\alpha J_{z}}=\cos(\alpha)J_{x}-\sin(\alpha)J_{y}.$ (51)
Subsequently, by using the Baker-Campbell-Hausdorff formula, we have
@@
where $C_{k+1}=igt_{1}[J_{z}\sigma_{z},C_{k}]$ and $C_{0}=J_{x}$. By using the commutation relation $[J_{\alpha},J_{\beta}]=i\epsilon_{\alpha\beta\gamma}J_{\gamma}$ with $\epsilon_{\alpha\beta\gamma}$ the Levi-Civita tensor, we have
@@
Appendix B QFI under parametric deviation
This appendix derives quantum Fisher information in the presence of deviations in eigen-frequency of the probe system $\omega_{P}$ and probe-ancillary qubit coupling strength $g$. By using the proper state of the ancillary qubit and the optimized joint-evolution time given in Eq. (18), QFI in Eq. (16) can be expressed as
$F_{Q}$ $=\sum_{i=1}^{d}4p_{i}\langle\cos^{2}(gt_{1,{\rm opt}})J^{2}(-\omega_{P}t_{1,{\rm opt}})\rangle_{\psi_{i}}$ (58)
$+\sum_{i=1}^{d}4p_{i}\left\langle\sin^{2}(gt_{1,{\rm opt}})J^{2}\left(\frac{\pi}{2}-\omega_{P}t_{1,{\rm opt}}\right)\right\rangle_{\psi_{i}}$
$-\sum_{i,j=1}^{d}\frac{8p_{i}p_{j}}{p_{i}+p_{j}}|\langle\psi_{i}|\cos(gt_{1,{\rm opt}})J(-\omega_{P}t_{1,{\rm opt}})|\psi_{j}\rangle|^{2}.$
When $g\rightarrow g^{\prime}=g+\Delta g$ and $\omega_{P}\rightarrow\omega_{P}^{\prime}=\omega_{P}+\Delta\omega_{P}$ with $|\Delta gt_{1,{\rm opt}}|\ll 1$ and $|\Delta\omega_{P}t_{1,{\rm opt}}|\ll 1$ the relative deviations, QFI for the optimized probe state $\rho_{P}^{\rm opt}=|j,j\rangle_{\rm opt}\langle j,j|$ is
$F_{Q}$ $\approx 4(1-\Delta g^{2}t_{1,{\rm opt}}^{2})$ (59)
×$\times_{\rm opt}\langle j,j|J^{2}\left[\frac{\pi}{2}-(\omega_{P}+\Delta\omega_{P})t_{1,{\rm opt}}\right]|j,j\rangle_{\rm opt}$
$+4\Delta g^{2}t_{1,{\rm opt}}^{2}\Big{\{}_{\rm opt}\langle j,j|J^{2}[-(\omega_{P}+\Delta\omega_{P})t_{1,{\rm opt}}]|j,j\rangle_{\rm opt}$
$-|_{\rm opt}\langle j,j|J[-(\omega_{P}+\Delta\omega_{P})t_{1,{\rm opt}}]|j,j\rangle_{\rm opt}|^{2}\Big{\}}.$
Here we use the approximations $\cos^{2}(g^{\prime}t_{1,{\rm opt}})\approx\Delta g^{2}t_{1,{\rm opt}}^{2}$ and $\sin^{2}(g^{\prime}t_{1,{\rm opt}})\approx 1-\Delta g^{2}t_{1,{\rm opt}}^{2}$. Using Taylor expansion, the three components of QFI in Eq. (59) can be approximated as
${}_{\rm opt}\langle j,j|J^{2}\left[\frac{\pi}{2}-(\omega_{P}+\Delta\omega_{P})t_{1,{\rm opt}}\right]|j,j\rangle_{\rm opt}$ (60)
$=$ $\langle j,j|e^{i\Delta\omega_{P}t_{1,{\rm opt}}J_{y}}J_{z}^{2}e^{-i\Delta\omega_{P}t_{1,{\rm opt}}J_{y}}|j,j\rangle$
$\approx$ $j^{2}-\frac{j(2j-1)}{2}\Delta\omega_{P}^{2}t_{1,{\rm opt}}^{2},$
${}_{\rm opt}\langle j,j|J^{2}[-(\omega_{P}+\Delta\omega_{P})t_{1,{\rm opt}}]|j,j\rangle_{\rm opt}$
$=$ $\langle j,j|e^{i\Delta\omega_{P}t_{1,{\rm opt}}J_{y}}J_{x}^{2}e^{-i\Delta\omega_{P}t_{1,{\rm opt}}J_{y}}|j,j\rangle$
$\approx$ $\frac{j}{2}+\frac{j(2j-1)}{2}\Delta\omega_{P}^{2}t_{1,{\rm opt}}^{2},$
${}_{\rm opt}\langle j,j|J[-(\omega_{P}+\Delta\omega_{P})t_{1,{\rm opt}}]|j,j\rangle_{\rm opt}$
$=$ $\langle j,j|e^{i\Delta\omega_{P}t_{1,{\rm opt}}J_{y}}J_{x}e^{-i\Delta\omega_{P}t_{1,{\rm opt}}J_{y}}|j,j\rangle$
$\approx$ $j\Delta\omega_{P}t_{1,{\rm opt}},$
respectively. Substituting Eq. (60) to Eq. (59), one can obtain QFI up to the quadratic terms about the relative deviations:
$F_{Q}\approx N^{2}-N(N-1)(\Delta\omega_{P}^{2}+\Delta g^{2})t_{1}^{2}.$ (61)
Then we obtain Eq. (22).